Agenda/Minutes of the VIS Meeting on 2020/7/3
2020/7/3 14:00 - 15:15
Participants: Takahashi, Hirata, Miyo, Sato, Fabian, Tanaka, Shoda
[Miyo] No progress about the heater jacket Issue
[Miyo] No progress about the preparation for new ACC test in KAGRA. Issue
- Preparing for the site test of the modified ACC. Travel for Kamioka is uncertain. The test using the DGS in TAMA/ATC was suggested.
- I checked Terrences' package Kontrol. I played with the optimizatiom of the blending filter:
- Type-B paper: I went back to the simulation of the injected control noise and found some wrong signs in my simulation. Now I'm beginning to get something similar to what Takanori got for the inertial damping, but it is still work in progress.
- PR models are updated to have control loop for PSD in MISALIGNED state.
At that time, PRs tripped due to the model restart failure klog.
- There was a trouble that the burt file was properly loaded, and optic was too away from the linear range to turn on the control without DC offset.
- Adjusted the actuation efficiency of TM and MN in the suspension models.
- Measuring the filters for the LVDT combiner. Small non-polarized electrolytic condensers for narrow space had no good combined capacitance. The film condensers are tested on the breadboard now.
- Post Obs. tasks are still on going. RT model modification for SRs was carried out. Sensor correction for SRs and SR suspension update are planed in the next week.
Schedule towards O4 (Chart) was revised.
- Takahashi (7/14-16)
- On 2020/7/10(Fri)